17

ADVANCED ANALYSIS Assume that demand for a commodity is represented by the equation P 20-20s Supply is represented by the equation where Qdand Qs are quantity demanded and quantity supplied, respectively, and Pis price Instructions: Round your answer for price to 2 decimal places and enter your quantity as a whole number. a. Using the equilibrium condition Qs Od, determine equilibrium price. b. Now determine equilibrium quantity. units
Answer #1

From the demand equation,

Qd = 10 - P/2

From the supply equation,

- Qs = -5/3 -P/3

Qs = 5/3 + P/3

At equilibrium, Qd = Qs

10 - P/2 = 5/3 + P/3

25/3 = 5P/6

5 = P/2

P = $10

Replacing P in either demand or supply equation, we get,

Q = 10 - 5 = 5 units